# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ 1.2 Related Work
Graph sampling methodologies are often propelled by many practical factors. Most notably, these encompass data scarcity, high data acquisition costs [SWM05], and limited surveys of hidden structures [LF09, YML13, FH16]. In scenarios where observations are sampled from two large networks, it becomes unrealistic to presume that correlation exists among all nodes within the sampled subgraphs. As a result, a pair of partially correlated graphs emerge naturally. While the precise number of correlated nodes may not be accessible, we often have some partial knowledge on the scale. For instance, when the observations are induced subgraphs of randomly selected nodes, the number of correlated nodes follows a hypergeometric distribution that concentrates around the mean value.
Besides the recent literature on the graph alignment problem, the correlation detection is another related topic. Given a pair of graphs, their correlation detection is formulated as a hypothesis testing problem, wherein the null hypothesis assumes independent random graphs, while the alternative assumes edge correlation under a latent permutation. [BCL+19] proposed a hypothesis testing model for correlated Erdős-Rényi graphs and provided a pseudo-polynomial time algorithm for detection under certain conditions on the edge connection probability and average degree. [WXY23] established the sharp threshold for dense Erdős-Rényi graphs and determined the threshold within a constant factor for sparse Erdős-Rényi graphs. [DD23a] derived the sharp threshold for sparse Erdős-Rényi graphs by analyzing the densest subgraph. Additionally, [MWXY21] proposed a polynomial time algorithm for detection by counting trees when the correlation coefficient exceeds a constant value. It is natural to ask whether the correlation can be detected when only a subsample from the graphs is collected. The probabilistic model is similar to the one present in the current paper, and we leave the exploration as our future work.
@@
Numerous algorithms have been developed for the recovery problem. For example, [YG13] analyzed the percolation graph matching algorithm, [BCL+19] analyzed the problem using subgraph matching techniques, and [MX20] obtained an algorithm for the seeded setting based on a delicate analysis of local neighborhoods. However, these algorithms may be computationally inefficient. There are several polynomial-time algorithms for recovery, catering to different regimes correlation coefficients $\rho$. These include works by [BES80, Bol82, DCKG19, GM20, DMWX21, MRT23, MWXY23, DL23, MS24]. For instance, [MWXY23] proposed a polynomial-time algorithm for recovery by counting chandeliers when the correlation coefficient $\rho>\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant introduced in [Ott48]. Additionally, [DL23] introduced an efficient iterative polynomial-time algorithm for sparse Erdős-Rényi graphs when the correlation coefficient is a constant.
@@
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, [MWXY21] proved that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal.
@@
The maximal overlap in the form of (4) is a test statistic which aims to identify the mapping that maximizes the edge correlation between two graphs. It is known that finding the maximal overlap is an instance of quadratic assignment problem (QAP) [PRW94], which is NP-hard to solve or to approximate [MMS10]. There are many studies aiming to detect or recover latent structures based on the maximal overlap statistics [CK16, CK17, BCL+19, MX20, DMWX21, WXY22, WXY23, HM23]. Finally, we mention that the recent work [DDG24] approximated the maximal overlap within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs.
@@
Many properties of the correlated Erdős-Rényi graphs model have been extensively investigated. However, the strong symmetry and tree-like structure inherent in this model distinguish it significantly from graph models encountered in practical applications. Therefore, it is crucial to explore more general graph models. One such model is inhomogeneous random graph model, where the edge connecting probability varies among edges in the graph [RS23, SPT23, DFW23]. Besides, geometric random graph model [WWXY22, BB23, SNL+23, GL24], planted cycle model [MWZ23, MWZ24], planted subhypergraph model [DMW23] and corrupt model [AH23] have also been subjects of recent studies.
@@ 1.3 Notations
For any $n\in\mathbb{N}$, let $[n]\triangleq\{1,2,...,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and for all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
For a given graph $G$, let $V(G)$ denote its vertex set and $E(G)$ denote its edge set. Let $\mathsf{{v}}(G)=|V(G)|$ denote the order of $G$ and $\mathsf{{e}}(G)=|E(G)|$ denote size of $G$. For a set $V$, let $\tbinom{V}{2}\triangleq\{\{x,y\}:x,y\in V,x\neq y\}$ denote the collection of all subsets of $V$ of cardinality two. We also write $uv$ to denote an edge $\{u,v\}$. The induce subgraph of $G$ over a vertex set $V$ is denoted by $G[V]$. Given an injective mapping of vertices $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$, the induced injective mapping of edges is defined as $\pi^{{\mathsf{E}}}:\binom{S}{2}\mapsto\binom{V(G_{2})}{2}$ as $\pi^{{\mathsf{E}}}(uv)=\pi(u)\pi(v)$ for $u,v\in S$. We also succinctly write $\pi(e)=\pi^{{\mathsf{E}}}(e)$ for an edge $e$ when the meaning is clear from the context.
@@ 2 Correlated functional digraph
A mapping from a set to itself can be graphically represented as functional digraph (see, e.g., [Wes21, Definition 1.3.3]). Here we extend the notion to a mapping with different domain and range sets, where the elements from the two sets are correlated. While our focus in this section is on the mapping between the edges in $G_{1}$ and $G_{2}$, the graphical representation can be easily extended to mappings between two arbitrary finite sets such as vertices.
We first provide an equivalent description of the estimator in (4). Given a domain subset $S\subseteq\binom{V(G_{1})}{2}$ and an injective function $\pi:S\mapsto\binom{V(G_{2})}{2}$, we define the intersection graph ${\mathcal{H}}_{\pi}$ as
$V({\mathcal{H}}_{\pi})=V(G_{1}),\quad e\in E({\mathcal{H}}_{\pi})\text{ if andonly if }e\in E(G_{1})\cap S\text{ and }\pi(e)\in E(G_{2}).$
The estimator (4) maximizes the size of the intersection graph $|E({\mathcal{H}}_{\pi})|$. More generally, in our analysis in Section 3, we need to count the number of edges present in some subset ${\mathcal{E}}\subseteq S$ given by
$|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|=\sum_{e\in{\mathcal{E}}}{\mathbf{1}_{\left\{{e\in E({\mathcal{H}}_{\pi})}\right\}}}=\sum_{e\in{\mathcal{E}}}{\mathbf{1}_{\left\{{e\in E(G_{1})}\right\}}}{\mathbf{1}_{\left\{{\pi(e)\in E(G_{2})}\right\}}}.$ (5)
Due to the correlation between the edges in $G_{1}$ and $G_{2}$, the counters ${\mathbf{1}_{\left\{{e\in E({\mathcal{H}}_{\pi})}\right\}}}$ are correlated random variables. The main idea is to decompose ${\mathcal{E}}$ into independent parts. Specially, the correlation is prescribed by the underlying mapping $\pi^{*}$ as illustrated in Figure 1, where the correlated edges are red dashed lines. To formally describe all correlation relationships, we introduce the correlated functional digraph of a mapping $\pi$ between a pair of graphs.
@@
Let $\pi^{*}:S^{*}\mapsto T^{*}$ be the underlying mapping between correlated elements. The correlated functional digraph of the function $\pi:S\mapsto T$ is constructed as follows. Let the vertex set be $S\cup S^{*}\cup T\cup T^{*}$. We first add every edge $e\mapsto\pi(e)$ for $e\in S$, and then merge each pair of nodes $(e,\pi^{*}(e))$ for $e\in S^{*}$ into one node.
@@
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes under $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where the self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 2. Let ${\mathcal{P}}$ and ${\mathcal{C}}$ denote the collections of subsets of ${\mathcal{E}}$ belonging to different connected paths and cycles, respectively. Note that the sets from ${\mathcal{P}}$ and ${\mathcal{C}}$ are disjoint. Consequently,
$|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|=\sum_{P\in{\mathcal{P}}}|P\cap E({\mathcal{H}}_{\pi})|+\sum_{C\in{\mathcal{C}}}|C\cap E({\mathcal{H}}_{\pi})|,$
@@
In our model, the edge correlations are assumed to be homogeneous, and hence the distribution of $|P\cap E({\mathcal{H}}_{\pi})|$ and $|C\cap E({\mathcal{H}}_{\pi})|$ only depends on the size of the component. Let $\kappa^{{\mathsf{P}}}_{\ell}(t)$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)$ denote the cumulant generating functions of $|P\cap E({\mathcal{H}}_{\pi})|$ and $|C\cap E({\mathcal{H}}_{\pi})|$ with $|P|=|C|=\ell$, respectively, and we have
$\log\mathbb{E}\left[e^{t|P\cap E({\mathcal{H}}_{\pi})|}\right]=\kappa^{{\mathsf{P}}}_{|P|}(t),\qquad\log\mathbb{E}\left[e^{t|C\cap E({\mathcal{H}}_{\pi})|}\right]=\kappa^{{\mathsf{C}}}_{|C|}(t).$
The lower-order cumulants can be promptly calculated. For instance,
$\kappa^{{\mathsf{C}}}_{1}(t)$ $=\log(1+p_{11}(e^{t}-1)),$ (6)
It is however essential to establish upper bounds for higher-order cumulants in terms of lower-order ones. To this end, we introduce the following lemma.
@@
For any $\rho>0$, $10<p<$ 1, and $t>0$,
$\kappa^{{\mathsf{P}}}_{1}(t)\leq\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t)\quad\text{and}\quad\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t),\quad\forall~{}\ell\geq 2.$
@@
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|}\right]\leq\frac{|{\mathcal{E}}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$ (8)
@@
The proof of Lemma 1 is deferred to Section A.1. The special case that both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], the correlation relationships under which can be characterized by a permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. However, in our case, the domain and range of $\pi$ and $\pi^{*}$ could be different and we need to deal with delicate correlations among the edges involving both cycles and paths by Lemma 1.
3 Recovery by maximizing the size of intersection graph
In this section, we prove the possibility results by analyzing the estimator $\hat{\pi}$ given in (4). By the optimality condition, it suffices to show that, for any $\pi^{*}\in\mathcal{S}_{n,m}$, we have $e({\mathcal{H}}_{\pi^{*}})$ exceeds $\max_{\pi:d(\pi,\pi^{*})>\tau}e({\mathcal{H}}_{\pi})$ with high probability when the underlying correlation is specified by $\pi^{*}$, where the thresholds $\tau=0$ and $\delta m$ are for exact and partial recoveries, respectively. In the following, we fix $\pi^{*}$ and provide a general recipe for the upper bound of $\mathbb{P}_{\pi^{*}}[d(\hat{\pi},\pi^{*})=k]$. The overall error probability follows from the summation over the desired range of $k$.

### v2

@@ 1.2 Related Work
When analyzing the properties of graphs, several challenges arise, such as limited data, high acquisition costs [SWM05], and incomplete knowledge of hidden structures [LF09, YML13, FH16]. Due to these challenges, graph sampling is a powerful approach for exploring graph structure. When data is sampled from two large networks, it is often unrealistic to assume full correlation among nodes in the sampled subgraphs. This naturally leads to partially correlated graphs. While the exact number of correlated nodes may be unknown, we usually have some estimate of the size. As a simplification, our model considers the case where the size is fixed. For the case where the size of correlated nodes is unknown, we leave it as our future work. See more details in Remark 4.
@@
Numerous algorithms have been developed for the recovery problem. For example, percolation graph matching algorithm [YG13], subgraph matching algorithm [BCL+19], and local neighborhoods based algorithm [MX20]. However, these algorithms may be computationally inefficient. There are several polynomial-time algorithms for recovery, catering to different regimes correlation coefficients $\rho$. These include works by [BES80, Bol82, DCKG19, GM20, DMWX21, MRT23, MWXY23, DL23, MS24]. For instance, a polynomial-time algorithm for recovery is proposed in [MWXY23] by counting chandeliers when the correlation coefficient $\rho>\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant introduced in [Ott48]. Additionally, there exists an efficient iterative polynomial-time algorithm for sparse Erdős-Rényi graphs when the correlation coefficient is a constant [DL23].
@@
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, it is shown in [MWXY24] that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, the recent work [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal.
@@
The maximal overlap is a test statistic which aims to identify the mapping that maximizes the edge correlation between two graphs. It is known that finding the maximal overlap is an instance of quadratic assignment problem (QAP) [PRW94], which is NP-hard to solve or to approximate [MMS10]. There are many studies aiming to detect or recover latent structures based on the maximal overlap statistics [CK16, CK17, BCL+19, MX20, DMWX21, WXY22, WXY23, HM23]. Finally, we mention that the recent work [DDG24] approximated the maximal overlap within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs. In this paper, we do not address polynomial algorithms or the computational hardness on partial models and leave these topics for future work.
Correlation detection
Besides the recent literature on the graph alignment problem, the correlation detection is another related topic. Given a pair of graphs, their correlation detection is formulated as a hypothesis testing problem, wherein the null hypothesis assumes independent random graphs, while the alternative assumes edge correlation under a latent permutation. A hypothesis testing model for correlated Erdős-Rényi graphs is proposed in [BCL+19]. The sharp threshold for dense Erdős-Rényi graphs and the threshold within a constant factor for sparse Erdős-Rényi graphs on this hypothesis testing model are established in [WXY23]. It is shown in [DD23a] that the sharp threshold for sparse Erdős-Rényi graphs can be derived by analyzing the densest subgraph. Additionally, a polynomial time algorithm for detection is also possible by counting trees when the correlation coefficient exceeds a constant value [MWXY24]. It is natural to ask whether the correlation can be detected when only a subsample from the graphs is collected. The probabilistic model is similar to the one present in the current paper, and we leave the exploration as our future work.
@@
Many properties of the correlated Erdős-Rényi model have been extensively investigated. However, the strong symmetry and tree-like structure inherent in this model distinguish it significantly from graph models encountered in practical applications. Therefore, it is crucial to explore more general graph models. One such model is inhomogeneous random graph model, where the edge connecting probability varies among edges in the graph [RS23, SPT23, DFW23]. Besides, geometric random graph model [WWXY22, SNL+23, BB24, GL24], planted cycle model [MWZ23, MWZ24], planted subhypergraph model [DMW23] and corrupt model [AH24] have also been subjects of recent studies.
@@ 1.3 Notations
For any $n\in\mathbb{N}$, let $[n]\triangleq\{1,2,\cdots,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
For a given weighted graph $G$, let $V(G)$ denote its vertex set and $E(G)$ its edge set. For a set $V$, let $\binom{V}{2}\triangleq\{\{x,y\}:x,y\in V,x\neq y\}$ denote the collection of all subsets of $V$ with cardinality two. We write $uv$ to represent an edge $\{u,v\}$, and $\beta_{e}(G)$ for the weight of the edge $e$. For unweighted graphs $G$, we define $\beta_{uv}(G)={\mathbf{1}_{\left\{{uv\in E(G)}\right\}}}$. Let $\mathsf{{v}}(G)=|V(G)|$ denote the number of vertices in $G$, and $\mathsf{{e}}(G)=\sum_{e\in E(G)}\beta_{e}(G)$ the total weight of its edges. The induced subgraph of $G$ over a vertex set $V$ is denoted by $G[V]$. Given an injective mapping of vertices $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$, the induced injective mapping of edges is defined as $\pi^{{\mathsf{E}}}:\binom{S}{2}\mapsto\binom{V(G_{2})}{2}$, where $\pi^{{\mathsf{E}}}(uv)=\pi(u)\pi(v)$ for $u,v\in S$. For simplicity, we write $\pi(e)$ to denote $\pi^{{\mathsf{E}}}(e)$ when it is clear from the context.
@@ 2 Correlated functional digraph
A mapping from a set to itself can be graphically represented as a functional digraph (see, e.g., [Wes21, Definition 1.3.3]). Here, we extend this notion to mappings between distinct domain and range sets, where elements from the two sets are correlated. Although our focus in this section is on the mappings between the edges of $G_{1}$ and $G_{2}$, the graphical representation can be easily generalized to mappings between any two finite sets, such as sets of vertices.
Given a domain subset $S\subseteq V(G_{1})$, an injective function $\pi:S\mapsto V(G_{2})$, and a bivariate function $f:\mathbb{R}\times\mathbb{R}\mapsto\mathbb{R}$, we define the $f$-intersected graph ${\mathcal{H}}_{\pi}^{f}$ as
$V({\mathcal{H}}_{\pi}^{f})=V(G_{1}),\quad\beta_{e}(\mathcal{H}_{\pi}^{f})=\begin{cases}f\left(\beta_{e}(G_{1}),\beta_{\pi^{\mathsf{E}}(e)}(G_{2})\right),&\text{ if }e\in E(G_{1})\cap\binom{S}{2},\\ 0,&\text{ otherwise.}\end{cases}$
The weight in ${\mathcal{H}}_{\pi}^{f}$ represents the similarity score under $\pi$. To establish the possibility results, our estimator maximizes the total similarity score:
$\hat{\pi}(f)$ $\in\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\mathsf{{e}}(\mathcal{H}_{\pi}^{f})=\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\sum_{e\in E(G_{1})}\beta_{e}(\mathcal{H}_{\pi}^{f}).$ (6)
Specifically, we use $\hat{\pi}(f)$ with $f(x,y)=xy$ to maximize overlap, and $f(x,y)=-\frac{1}{2}(x-y)^{2}$ to minimize the mean-squared distance. For notational simplicity, we write $\hat{\pi}=\hat{\pi}(f)$ when the choice of $f$ is clear from the context.
More generally, in our analysis in Section 3, we need to calculate the total weight within a subset ${\mathcal{E}}\subseteq\binom{S}{2}$ given by
$\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})\triangleq\sum_{e\in{\mathcal{E}}}\beta_{e}(\mathcal{H}_{\pi}^{f}).$ (7)
Due to the correlation between the edges in $G_{1}$ and $G_{2}$, the summands $\beta_{e}(\mathcal{H}_{\pi}^{f})$ are correlated random variables. The main idea is to decompose ${\mathcal{E}}$ into independent parts. Specially, the correlation is governed by the underlying mapping $\pi^{*}$, as illustrated in Figure 1 where the correlated edges are represented by red dashed lines. To formally describe all correlation relationships, we introduce the correlated functional digraph of a mapping $\pi$ between a pair of graphs.
@@
Let $\pi^{*}:S^{*}\mapsto T^{*}$ be the underlying mapping between correlated elements. The correlated functional digraph of the function $\pi:S\mapsto T$ is constructed as follows. Let the vertex set be $\binom{S}{2}\cup\binom{S^{*}}{2}\cup\binom{T}{2}\cup\binom{T^{*}}{2}$. We first add every edge $e\mapsto\pi(e)$ for $e\in\binom{S}{2}$, and then merge each pair of nodes $(e,\pi^{*}(e))$ for $e\in\binom{S^{*}}{2}$ into one node.
@@
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes according to $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where a self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 2. Let ${\mathcal{P}}$ and ${\mathcal{C}}$ denote the collections of subsets of ${\mathcal{E}}$ belonging to different connected paths and cycles, respectively. Note that the sets from ${\mathcal{P}}$ and ${\mathcal{C}}$ are disjoint. Consequently,
$\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})=\sum_{P\in{\mathcal{P}}}\beta_{P}(\mathcal{H}_{\pi}^{f})+\sum_{C\in{\mathcal{C}}}\beta_{C}(\mathcal{H}_{\pi}^{f}),$
@@
In our models, edge correlations are assumed to be homogeneous, implying that the distributions of $\beta_{P}(\mathcal{H}_{\pi}^{f})$ and $\beta_{C}(\mathcal{H}_{\pi}^{f})$ depend only on the size of the component. Let $\kappa^{{\mathsf{P}},f}_{\ell}(t)$ and $\kappa^{{\mathsf{C}},f}_{\ell}(t)$ denote the cumulant generating functions of $\beta_{P}(\mathcal{H}_{\pi}^{f})$ and $\beta_{C}(\mathcal{H}_{\pi}^{f})$ for components of order $\ell$, respectively. Then, we have
$\log\mathbb{E}\left[e^{t\beta_{P}(\mathcal{H}_{\pi}^{f})}\right]=\kappa^{{\mathsf{P}},f}_{|P|}(t),\qquad\log\mathbb{E}\left[e^{t\beta_{C}(\mathcal{H}_{\pi}^{f})}\right]=\kappa^{{\mathsf{C}},f}_{|C|}(t).$
For simplicity, we write $\kappa_{\ell}^{\mathsf{C}}(t)=\kappa_{\ell}^{{\mathsf{C}},f}(t)$ and $\kappa_{\ell}^{\mathsf{P}}(t)=\kappa_{\ell}^{{\mathsf{P}},f}(t)$ when the function $f$ is specified. Lower-order cumulants can be calculated directly. See more details in Appendix B.1. However, it is crucial to establish upper bounds for higher-order cumulants in terms of the lower-order ones. To this end, we introduce the following lemma.
@@
For any $10<p,\rho<$ 1,
$\kappa^{{\mathsf{P}}}_{1}(t)\leq\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t)\quad\text{and}\quad\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t),\quad\forall~{}\ell\geq 2,$
under any of the following three conditions:
  • In the Erdős-Rényi model with $f(x,y)=xy$ and $t>0$;
  • In the Gaussian Wigner model with $f(x,y)=xy$ and $0<t<\frac{1}{1+\rho}$;
  • In the Gaussian Wigner model with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ and $t>0$.
@@
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|{\mathcal{E}}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$ (8)
@@
The proof of Lemma 1 is deferred to Appendix B.1. The special case where both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], where the correlation relationships can be characterized by the permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. In contrast, in our setting, the domains and ranges of $\pi$ and $\pi^{*}$ may differ, requiring us to deal with the intricate correlations among edges involving both cycles and paths, as addressed by Lemma 1.
Remark 2.
For the Gaussian Wigner model, the MLE is defined by (6) with the following similarity function:
$f(x,y)=-\frac{\rho^{2}}{2}\left(x^{2}+y^{2}\right)+\rho xy.$ (9)
We analyze approximations of the above similarity score under different regimes. Specifically, when $\rho=1-\Omega(1)$, we apply the estimator with $f(x,y)=xy$, whereas for $\rho=1-o(1)$ the estimator with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ turns out to be crucial.
3 Recovery by maximizing total similarity score
In this section, we establish the possibility results by analyzing the estimators $\hat{\pi}$ defined in (6). For any two injections $\pi:S\mapsto T$ and $\pi^{\prime}:S^{\prime}\mapsto T^{\prime}$ with $|S|=|S^{\prime}|=m$, let $d(\pi,\pi^{\prime})\triangleq m-|\{v\in S\cap S^{\prime}:\pi(v)=\pi^{\prime}(v)\}|$. Indeed, $d(\pi,\pi^{\prime})=m(1-\mathrm{overlap}(\pi,\pi^{\prime}))$, where $\mathrm{overlap}(\pi,\pi^{\prime})$ is defined in (3). By the optimality condition, it suffices to show that
$\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})>\max_{\pi:d(\pi,\pi^{*})\geq d_{0}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})=\max_{k\geq d_{0}}\max_{\pi:d(\pi,\pi^{*})=k}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})$
with high probability, where the thresholds $d_{0}=1$ and $(1-\delta)m$ correspond to exact and partial recoveries, respectively. In the following, we outline a general recipe to derive an upper bound for the failure event $\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\}$ for a fixed $k$, where $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denotes the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. The error probabilities in the main theorems are then obtained by summing over the corresponding range of $k$.